OpenStudy (anonymous):

Given the equation y − 4 =3over4 (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

Is it?? \[y - 4 = \frac{3}{4}(x+8)\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Subtract 4 from both the sides and tell me what you get??

OpenStudy (anonymous):

y =3over4 x + 12

OpenStudy (anonymous):

No little bit wrong.. \[y = \frac{3}{4}x + 2 -1 \implies y = \frac{3}{4}x + 1\]

OpenStudy (anonymous):

so what is da final answer

OpenStudy (anonymous):

I have written it above...

OpenStudy (anonymous):

Slope Intercept form is given by: \[\huge \color{green}{y = mx + c}\]

OpenStudy (anonymous):

thanks but can post da final answer for me again

OpenStudy (anonymous):

What I have written after it \(\implies\) ?? That is the answer..

OpenStudy (anonymous):

THANKS BRO APPRECIATE IT